OpenStudy (anonymous):

Find the slope and y intercept of the line that passes through (2,8)(10,20)

hartnn (hartnn):

Use the equation y=mx+c The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and first find m.

OpenStudy (anonymous):

Ok so I got 1.5 for m

hartnn (hartnn):

correct! m= 1.5 or 3/2 now, you have y= 1.5x+c and a point on it , say lets take 2,8 so plug in x=2,y=8 in y=1.5x+c and find c= ... which is your y-intercept.

OpenStudy (anonymous):

c=2.6

OpenStudy (anonymous):

So how do I get my y intercept

OpenStudy (anonymous):

So my slope is 1.5 and y intercept 2.6

hartnn (hartnn):

plug in x=2,y=8 in y=1.5x+c 8= 1.5(2)+c ----> c= 8-3=5 c=5= y-intercept.

OpenStudy (anonymous):

ok so slope 1.5 and y intercept is 5

hartnn (hartnn):

yes, correct.

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^